OpenStudy (anonymous):

sqrt[5]{-\left(\begin{matrix}1 \\ 243\end{matrix}\right)} Simplify

OpenStudy (anonymous):

\[\sqrt[5]{-\left(\begin{matrix}1 \\ 243\end{matrix}\right)}\]

OpenStudy (anonymous):

-1/243 should be a fraction sorry it did not show up

OpenStudy (anonymous):

= -1 / 5th root of 243

OpenStudy (anonymous):

use ur calculator

OpenStudy (anonymous):

\[-\left(\begin{matrix}1 \\ 5\sqrt{243}\end{matrix}\right)\]

OpenStudy (anonymous):

My calculator tells me error. P.O.crap anyway, lol

jimthompson5910 (jim_thompson5910):

\[\sqrt[5]{\frac{-1}{243}}\] \[\sqrt[5]{\frac{(-1)^5}{(3^5)}}\] \[\sqrt[5]{\left(-\frac{1}{3}\right)^5}\] \[-\frac{1}{3}\] So \[\sqrt[5]{\frac{-1}{243}}=-\frac{1}{3}\] You're getting an error probably because some calculators don't like negatives in roots

OpenStudy (anonymous):

Oh ok, is it stil -1/3 if the whole fraction is negative?

OpenStudy (anonymous):

take 5th root of 243 and place - in front

OpenStudy (anonymous):

That's probably a dumb question sorry

jimthompson5910 (jim_thompson5910):

yes the answer is -1/3 because the fraction is negative.

OpenStudy (anonymous):

thats ok yes the answer is -1/3

jimthompson5910 (jim_thompson5910):

no not a dumb question at all

OpenStudy (anonymous):

Thank you both, I feel really dumb that I can not figure this stuff out